OpenStudy (anonymous):

Find the first four terms of the arithmetic sequence when a1=3 and=2. a. 2,5,8,11 b. 3,5,7,9 c. 1,3,5,7 d. 3,6,12,24

OpenStudy (saifoo.khan):

what's 2?

OpenStudy (anonymous):

d=2

OpenStudy (saifoo.khan):

1+2 = 1+2 + 2 = 1+2 + 2 + 2= 1 + 2 + 2 + 2 =

OpenStudy (anonymous):

Thank you!